OpenStudy (anonymous):

what polynomial has a graph that passes through the given points? (-4,-69),(-3,-28),(1,-4),(3,8)

OpenStudy (anonymous):

this is a long involved process solving a 4 by 4 system of equations takes for ever i would cheat http://www.wolframalpha.com/input/?i=polynomial+%28-4%2C-69%29%2C%28-3%2C-28%29%2C%281%2C-4%29%2C%283%2C8%29

OpenStudy (anonymous):

Lol I did the exact same thing for this one..beat me to it.

OpenStudy (anonymous):

thank you for the website

OpenStudy (anonymous):

wish they had this when i was in school i'd have 4 degrees by now

OpenStudy (whpalmer4):

It's not really so bad… \[y = ax^3+bx^2+cx+d\] \[-69 = a(-4)^3+b(-4)^2+c(-4)+d\]\[-69 = -64a+16b-4c+d\] \[-28=a (-3)^3+b(-3)^2+c(-3)+d\]\[-28 = -27a+9b-3c+d\] \[-4=a(1)^3+b(1)^2+c(1)+d\]\[-4=a+b+c+d\] \[8 = a(3)^3+b(3)^2+c(3)+d\]\[8=27a+9b+3c+d\] \[-69 = -64a+16b-4c+d\]\[-28 = -27a+9b-3c+d\]\[-4=a+b+c+d\]\[8=27a+9b+3c+d\] Use an equation solver to solve the system, and out pops \(a=1,b=-1,c=-3,d=-1\) You ought to at least understand how to do it, even if you're going to use some other tool most of the time!